# Beyond Incompatibility: Trade-offs between Mutually Exclusive Fairness Criteria in Machine Learning and Law

### v5

@@ 3.2 Mathematical Preliminaries on Optimal Transport
Our goal is to design an algorithm that interpolates smoothly between the three fairness criteria discussed, as they are typically not attainable at the same time. The interpolation relies on the mathematical theory of optimal transport, which we now wish to recall.
@@
where $T^{-1}(a,b)$ denotes the preimage of $(a,b)$ under $T$, i.e., the set of all numbers that $T$ maps to the interval $(a,b)$. One often uses the notation $\mu=\nu\circ T^{-1}$ in this situation. This means the following: If a random variable $X$ with values in $[0,1]$ is distributed according to $\nu$, then $T$ is a transport map from $\nu$ to $\mu$ if and only if the random variable $T(X)$ is distributed according to $\mu$. As an example on $\mathbb{R}$, consider the two normal distributions $\mathcal{N}(0,1)$ and $\mathcal{N}(1,2)$. Then $T:\mathbb{R}\to\mathbb{R}$, $T(x)=2x+1$, would be a transport map from $\mathcal{N}(0,1)$ to $\mathcal{N}(1,2)$. However, there may be many transport maps: In this example, also $x\mapsto-2x+1$ would be a transport map. Among the many transport maps, one (or potentially several) may be optimal in the following sense:
@@
Let $\nu,\mu$ be two absolutely probability measures on $[0,1]$ with finite variance. An optimal transport map is a transport map between $\nu$ and $\mu$ that minimises the cost functional
@@
Clearly, $\nu^{0}=\nu$ and $\nu^{1}=\mu$. The measure $\nu^{\theta}$ thus is a measure ‘in between’ $\nu$ and $\mu$ that is closer to $\nu$ the smaller $\theta$ is chosen and closer to $\mu$ the larger it is.
@@
$\{\nu_{k}\}_{1,\ldots,N}$ be a family of absolutely continuous probability with finite variance, and let $\{w_{k}\}_{k=1,\ldots,N}$ be positive weights with $\sum_{k=1}^{N}w_{k}=1$. Then there exists a unique probability measure $\nu$ on $[0,1]$ that minimizes the functional
@@
With two groups and respective score distributions $\nu_{1},\nu_{2}$ and weights $w_{1},w_{2}$, the barycenter is given by the displacement interpolation $\nu=\nu_{1}\circ(T^{w_{2}})^{-1}$, where $T$ is the optimal transport map between $\nu_{1}$ and $\nu_{2}$ and $T^{w_{1}}$ is given in Def.  4.
@@ 32       $\bar{\mu}_{t}$ = wasserstein2barycenter($\mu^{A}_{t},\mu^{B}_{t},\mu^{C}_{t}$, thetas[$t$])
We do this as follows (see again Algorithm 1): First (Lines 11), we compute said conditional probabilities $\lambda_{t}^{-}$ and $\nu_{t}$ from the given data. By Bayes’ Theorem, the probability distribution for the score of an individual $x$ from group $t$ conditional on $x$ being negative is given as
@@
Secondly, we consider the Wasserstein-2 barycenter of $\sigma^{-}_{1},\sigma^{-}_{2}$ on $[0,1]$ with weights $\frac{N_{1}}{N_{1}+N_{2}}$ and $\frac{N_{2}}{N_{1}+N_{2}}$, respectively, which we denote by $\bar{\sigma}^{-}$ (Line 1). Then there are two (optimal) transport maps $T^{-}_{t}:[0,1]\to[0,1]$ such that
@@ 3.3.4 Combining the three procedures
Each of the modified score maps $S_{t}^{A},S_{t}^{B},S_{t}^{C}$ gives rise to a corresponding score distribution $\mu_{t}^{A},\mu_{t}^{B},\mu_{t}^{C}$. For each group $t=1,2$, we therefore obtain a triangle in Wasserstein space. Let now $\theta_{t}^{A},\theta_{t}^{B},\theta_{t}^{C}\in[0,1]$ be three real parameters that add up to $1$, which are given as input to the algorithm. We then consider the weighted barycenter $\bar{\mu}_{t}$ in Wasserstein-2 space, i.e., the unique probability measure $\bar{\mu}$ that minimizes the functional
@@
A final remark: By construction, whenever we optimize for only one of the fairness criteria (that is, one of the $\theta$’s is set to one, this respective criterion will be exactly satisfied, whereas the other two might actually get worse. An interesting mathematical question is whether, when the three criteria are ‘mixed’ (i.e., none of the $\theta$’s is zero), it is possible that all three fairness goals simultaneously deteriorate – a sort of lose-lose situation. While we cannot exclude that this happens in pathological cases, it is certainly not an expected effect in realistic circumstances, and we have not observed such a lose-lose scenario in our experiments to be discussed in the next section.
@@ 4.3.1 Data set
For our last experiments, we use data from the e-commerce platform Zalando, one of Europe’s largest fashion and lifestyle retailers, operating in 25 European countries. We collected a data set of 81 048 articles on 27th October 2021, which belong to four different women clothes categories with a similar price range: skirts, jeans, trousers, and knitwear. Each data point contains the following information: a brand, a ranking score, and the number of impressions and clicks during the week after article collection. Products are labeled as relevant (ground truth positive) if the ratio of clicks to impressions is above a certain threshold. All other products are labeled not relevant (ground truth negative). Groups are assigned as follows: we compute a brand’s visibility as the average number of impressions per article for each brand and each category, and thus categorize brands having either low or high visibility. When this average is below the median, a brand is said to belong to the ‘low’ group, while if the number is above the third quartile, a brand is set to belong to the ‘high’ group. Articles of brands with visibility between the low and high group thresholds are discarded, leaving us with 62 461 articles. 70.5% of them belong to high visibility brands. Note, that the ranking scores are produced by a learning-to-rank model trained daily to optimize a surrogate normalized Discounted Cumulative Gain (nDCG) loss [61], but not to classify articles according to the ground truth labeling we defined above. However, because the nDCG relevance labels are based on customer interactions (i.e., implicit feedback), it is reasonable to assume that a good ranking model should be able to identify which article is likely to be clicked. Hence, the scores of an accurate click-through-rate classifier should provide a decent ranking. We observe in our data that, irrespective of whether high visibility articles are relevant or not, they receive higher ranking scores compared to low visibility articles. There are many reasons external to the ranking algorithm why customers would prefer well-known and highly visible brands: these brands have higher marketing budgets, they can afford a more aggressive pricing strategy thanks to economies of scale, and they benefit from the Matthew effect of accumulated advantage [84]. Moreover, it is in the commercial interest of the platform to highlight such best-selling items. At the same time, the platform may not want to reinforce the status quo but rather provide a level playing field for all brands to compete fairly. A level playing field makes it easier to attract new brands on the platform, offers a more diverse assortment to a potentially larger customer base, and can be seen as a good step to address the legal requirements discussed in 5.3.
@@ 4.3.2 Results
Figure 4: (Best seen in color). Product rankings with low and high visibility brands under different settings. Colors indicate group membership, stripes indicate a positive label, i.e., these are the products customers find more relevant. Fig. 4a exemplifies an optimal outcome, where all relevant products are in the top positions, and the visibility of brands A and B is distributed according to their share of relevant products. Fig. 4b shows the distribution of relevant (and non-relevant) products from low and high visibility brands as shown by Zalando’s ranking algorithm by the time of data collection. We see that famous brands do indeed have an advantage over less-known brands, as there are almost no products from the low group in the first few bins. Fig. 4c–f show how the ranking changes after FAIM is applied with different values for $\theta^{A}$, $\theta^{B}$, and $\theta^{C}$. Depending on our preference for calibration, balance for the positive class, or balance for the negative class, we see varying degrees of visibility improvement for the low group.
@@
From Table 3 we see that low visibility brands do indeed have a relative disadvantage over high visibility brands. The low recall, and the high false negative rate indicate that many relevant items from the low group are not shown to the customer. This situation is improved by FAIM in all four scenarios, albeit to varying degrees. Our results also show that the algorithm yields expected improvements: When criterion A) is desired, accuracy and recall is improved for both groups. Pursuing criterion B) or C) levels the respective error rates. Finally, a compromise between all three criteria is found when setting $\theta^{A}=\theta^{B}=\theta^{C}$.
@@
For a better understanding on the effect of FAIM on the rankings, observe Fig. 4. In Fig. 4a we show a fictive example for an ideal ranking: imagine a data set with two groups A and B. A has 20,000 items: 4,000 positive and 16,000 negative. B has 10,000 items: 6,000 positive and 4,000 negative. If we assume a perfect ranker, we can sort those items simply by their score. When we group them in bins of 2 000 items each (to be consistent with how Zalando articles are binned), we would expect the first 10 bins to contain only positive (relevant) items, and the next 20 bins to have only negative (non-relevant) items. Furthermore, the positive bins should display a ratio of 6000/4000 for groups B and A respectively, while the negative bins should display a ratio of 4000/16000. Fig. 4b shows the ranking as collected from the Zalando website on 27th October 2021, which by virtue of being real is not ideal. Since each bin contains more than 2 000 items, it is enough to focus on the first bins, as other items are shown very rarely, unless customers are actively looking for them. We see that the first five bins contain mostly high visibility brands, meaning that relevant products from low visibility brands are rarely shown to the customer, unless they start to use product filters or full text search. Again, there are various reasons why a customer would prefer well-known over less-known brands, but if we wanted to (or had to because of legislation) mitigate the visibility disparities between high and low brands, FAIM provides a convenient approach (see Section 5.3 for a discussion on FAIM’s legal meaning for an e-commerce scenario such as ours). Figures 4c– 4f show rankings after FAIM has been applied with different values for $\theta^{A}$, $\theta^{B}$, and $\theta^{C}$. In all cases, FAIM distributes visibility better across both brand groups, particularly for relevant products, which is the interesting case for an e-commerce platform such as Zalando.
@@ 5.1 Recidivism Prediction
In this paper, we consider recidivism prediction for technical and argumentative reasons: the original papers by Kleinberg et al. [67] and Chouldechova [20]showing incompatibility between fairness metrics did use the COMPAS case, and we directly build on their work. Hence, we use COMPAS as an example to clarify how FAIM works. However, when considering recidivism prediction, the first questions that arise from a juridical perspective are whether one legally may, or policy-wise should, use algorithmic tools such as the COMPAS model to assess recidivism risk in Criminal Justice proceedings at all (see also 6.3). We do not wish to make any claim as to the legitimacy of the usage of these tools in this paper; there are, if anything, good reasons to be quite skeptical about it.
@@
As mentioned, COMPAS is a proprietary model used in pretrial settings in the US to determine whether potential offenders should be detained until their trial, based on a prediction of recidivism likelihood. Similar models are currently employed in Canada, the UK, and Spain [107, p. 122]. While the use of such instruments has sparked significant controversy, and critique, in the legal literature [98, 76, 93, 64, 28, 59, 89], courts such as the Wisconsin Supreme Court have condoned their use under certain safeguards,888 State v. Loomis, 881 N.W.2d 749 (Wis. 2016). and their deployment is currently on the rise [52]; [107, p. 122].
Within the scope of this paper, we cannot offer an in-depth discussion of the promises and perils of algorithmic recidivism risk assessment (see, e.g., [96, 68], and 6.3). Rather, we would like to point out that, to the extent that such tools are used at all, they must clearly fulfill minimum requirements seeking to safeguard normative and legal principles of the jurisdictions in which they are deployed. Importantly, the impossibility theorems concerning various fairness metrics mentioned above apply equally in the case of algorithmic and human risk assessments. While the influence of fairness considerations on human risk predictions, for example by a judge, remains difficult to elucidate, FAIM allows to make the relevant trade-offs transparent and to rank the involved fairness metrics in varying degrees of priority.

### v6

@@ 3.2 Mathematical Preliminaries on Optimal Transport
Our goal is to design an algorithm that interpolates smoothly between the three fairness criteria discussed, as they are typically not attainable at the same time, as discussed in the preceding subsection. The interpolation relies on the mathematical theory of optimal transport, which we now wish to recall.
@@
where $T^{-1}(a,b)$ denotes the preimage of $(a,b)$ under $T$, i.e., the set of all numbers that $T$ maps to the interval $(a,b)$. One often uses the notation $\mu=\nu\circ T^{-1}$ in this situation. This means the following: If a random variable $X$ with values in $[0,1]$ is distributed according to $\nu$, then $T$ is a transport map from $\nu$ to $\mu$ if and only if the random variable $T(X)$ is distributed according to $\mu$. As an example on $\mathbb{R}$, consider the two normal distributions $\mathcal{N}(0,1)$ and $\mathcal{N}(1,4)$, represented by Gaussian bell functions with expectation 0 and variance 1, or with expectation 1 and variance 4, respectively. Then $T:\mathbb{R}\to\mathbb{R}$, $T(x)=2x+1$, would be a transport map from $\mathcal{N}(0,1)$ to $\mathcal{N}(1,4)$. However, there may be many transport maps: In this example, also $x\mapsto-2x+1$ would be a transport map. Among the many transport maps, one (or potentially several) may be optimal in the following sense:
@@
Let $\nu,\mu$ be two absolutely probability measures on $[0,1]$. An optimal transport map is a transport map between $\nu$ and $\mu$ that minimises the cost functional
@@
Clearly, $\nu^{0}=\nu$ and $\nu^{1}=\mu$. The measure $\nu^{\theta}$ thus is a measure ‘in between’ $\nu$ and $\mu$ that is closer to $\nu$ the smaller $\theta$ is chosen and closer to $\mu$ the larger it is. As an example, consider on the real numbers the normal distributions $\nu=\mathcal{N}(0,1)$ and $\mu=\mathcal{N}(1,1)$. The optimal transport maps from $\nu$ to $\mu$ in this case is $T(x)=x+1$, so that $T^{\theta}=x+\theta$, and correspondingly $\nu^{\theta}=\mathcal{N}(\theta,1)$. Note this is not the same as the convex combination of the distributions, which would be $(1-\theta)\mathcal{N}(0,1)+\theta\mathcal{N}(1,1)$.
@@
$\{\nu_{k}\}_{1,\ldots,N}$ be a family of absolutely continuous probability distributions, and let $\{w_{k}\}_{k=1,\ldots,N}$ be positive weights with $\sum_{k=1}^{N}w_{k}=1$. Then there exists a unique probability measure $\nu$ on $[0,1]$ that minimizes the functional
@@
Thus, the barycenter gives the least square minimization on the level of probability measures, so to speak. With two groups and respective score distributions $\nu_{1},\nu_{2}$ and weights $w_{1},w_{2}$, the barycenter is given by the displacement interpolation $\nu=\nu_{1}\circ(T^{w_{2}})^{-1}$, where $T$ is the optimal transport map between $\nu_{1}$ and $\nu_{2}$ and $T^{w_{2}}$ is given in Def. 4. Specifically, in our example with $\nu=\mathcal{N}(0,1)$ and $\mu=\mathcal{N}(1,1)$ and weights, say, $w_{1}=0.2$ and $w_{2}=0.8$, the barycenter would be $\mathcal{N}(0.8,1)$.
@@ 32       $\bar{\mu}_{t}$ = wasserstein2barycenter($\mu^{A}_{t},\mu^{B}_{t},\mu^{C}_{t}$, thetas[$t$])
We do this as follows (see again Algorithm 1): First (Lines 11), we compute said conditional probabilities $\lambda_{t}^{-}$ and $\nu_{t}$ from the given data. By Bayes’ Theorem for continuous random variables, the probability distribution for the score of an individual $x$ from group $t$ conditional on $x$ being negative is given as
@@
Secondly, we consider the Wasserstein-2 barycenter of $\sigma^{-}_{1},\sigma^{-}_{2}$ on $[0,1]$ with weights $\frac{N_{1}}{N_{1}+N_{2}}$ and $\frac{N_{2}}{N_{1}+N_{2}}$, respectively, which we denote by $\bar{\sigma}^{-}$ (Line 1). Then there are two (optimal) transport maps $T^{-}_{t}:[0,1]\to[0,1]$ between $\sigma^{-}_{1}$ and $\bar{\sigma}^{-}$, and between $\sigma^{-}_{2}$ and $\bar{\sigma}^{-}$, such that
@@ 3.3.4 Combining the three procedures
Each of the modified score maps $S_{t}^{A},S_{t}^{B},S_{t}^{C}$ gives rise to a corresponding score distribution $\mu_{t}^{A},\mu_{t}^{B},\mu_{t}^{C}$. For each group $t=1,2$, we therefore obtain a triangle in Wasserstein space, i.e., in the space of probability measures. Let now $\theta_{t}^{A},\theta_{t}^{B},\theta_{t}^{C}\in[0,1]$ be three real parameters that add up to $1$, which are given as input to the algorithm. We then consider the weighted barycenter $\bar{\mu}_{t}$ in Wasserstein-2 space, i.e., the unique probability measure $\bar{\mu}$ that minimizes the functional
@@
A final remark: By construction, whenever we optimize for only one of the fairness criteria (that is, one of the $\theta$’s is set to one), this respective criterion will be exactly satisfied, whereas the other two might actually get worse. An interesting mathematical question is whether, when the three criteria are ‘mixed’ (i.e., none of the $\theta$’s is zero), it is possible that all three fairness goals simultaneously deteriorate – a sort of lose-lose situation. While we cannot exclude that this happens in pathological cases, it is certainly not an expected effect in realistic circumstances, and we have not observed such a lose-lose scenario in our experiments to be discussed in the next section.
@@ 4.3.1 Data set
For our last experiments, we use data from the e-commerce platform Zalando, one of Europe’s largest fashion and lifestyle retailers, operating in 25 European countries. We collected a data set of 81 048 products for-sale on 27th October 2021, which belong to four different women clothes categories with a similar price range: skirts, jeans, trousers, and knitwear. Each data point contains the following information: a brand, a ranking score, number of impressions (how often the product was seen), and clicks. Impressions and clicks correspond to the week following 27th October, 2021. Products are labeled as relevant (ground truth positive) if the ratio of clicks to impressions is above a certain threshold, meaning that they were of interest to users when seen by the customer. All other products are labeled not relevant (ground truth negative). Groups are assigned based on brand visibility calculated as the average number of impressions over all products per brand per category. When this average is below the median, a brand is said to belong to the ‘low’ brand visibility group. When this average is above the third quartile, a brand is said to belong to the ‘high’ brand visibility group. Products of brands with visibility between the low and high group thresholds are discarded, leaving 62 461 products. 70.5% of them belong to high visibility brands. Note that the ranking scores are produced by a learning-to-rank model trained daily to optimize a surrogate normalized Discounted Cumulative Gain (nDCG) loss [67], but not to classify products according to the ground truth labeling we defined above. However, because the nDCG relevance labels are based on customer interactions (i.e., implicit feedback), it is reasonable to assume that a good ranking model should be able to identify which product is likely to be clicked. Hence, the scores of an accurate click-through-rate classifier should provide a decent ranking. We observe in our data that, irrespective of whether high visibility products are relevant or not, they receive higher ranking scores compared to low visibility products. There are many reasons external to the ranking algorithm why customers would prefer well-known and highly visible brands: these brands have higher marketing budgets, they can afford a more aggressive pricing strategy thanks to economies of scale, and they benefit from the Matthew effect of accumulated advantage [92]. Moreover, it is in the commercial interest of the platform to highlight such best-selling items. At the same time, the platform may not want to reinforce the status quo but rather provide a level playing field for all brands to compete fairly. A level playing field makes it easier to attract new brands on the platform, offers a more diverse assortment to a potentially larger customer base, and can be seen as a good step to address the legal requirements regarding competition law and the Digital Markets Act discussed in section 5.3.
@@ 4.3.2 Results
Figure 4: (Best seen in color). Product rankings with low and high visibility brands under different settings. Color indicates the average overall brand visibility of a product (group membership: low and high), whereas fill pattern indicates customer relevance (ground truth: relevant and not relevant). Fig. 4a exemplifies the outcome of an ideal ranker that ranks products by relevance irrespective of brand (and brand visibility). Fig. 4b shows the distribution of relevant vs not relevent products from low and high visibility brands as shown by Zalando’s ranking algorithm by the time of data collection. We see that high visibility brands do indeed have an advantage over low visibility brands, as there are almost no products from the low brand visibility group in the first few bins. Fig. 4c–f show how the ranking changes after FAIM is applied with different values for $\theta^{A}$, $\theta^{B}$, and $\theta^{C}$. Depending on our preference for calibration, balance for the positive class, or balance for the negative class, we see varying degrees of ranking improvement for relevant products from the low brand visibility group.
@@
From Table 3 we see that low visibility brands do indeed have a relative disadvantage over high visibility brands. The low recall, and the high false negative rate indicate that many relevant products from the low group are not shown to the customer. This situation is improved by FAIM in all four scenarios, albeit to varying degrees. Our results also show that the algorithm yields expected improvements: When criterion A) is desired, accuracy and recall is improved for both groups. Pursuing criterion B) or C) levels the respective error rates. Finally, a compromise between all three criteria is found when setting $\theta^{A}=\theta^{B}=\theta^{C}$.
@@
For a better understanding on the effect of FAIM on the rankings, observe Fig. 4. In Fig. 4a we show a fictive example for an ideal ranking: imagine a data set with brand visibility groups. The low brand visibility group has 400 products: 200 relevant and 200 non relevant. The high brand visibility group has 1100 products: 300 relevant and 800 non relevant. If we assume a ideal ranker, we can sort those products simply by their relevance score. When grouping the rankings in bins of 50 products, we would expect the top 10 bins (top 500 ranked products) to contain only relevant products, and the next 20 bins (next 1000 ranked products) to contain only non relevant products. The top 10 (relevant) bins should contain a high to low brand visibility ratio of 300/200, whereas the next 20 (non relevant) bins should contain a high to low brand visibility ratio of 800/200. Fig. 4b shows the ranking as collected from the Zalando website on 27th October 2021, which, by virtue of being real, is not ideal. Since each bin contains more than 2 000 products, it is enough to focus on the first bins, as products from other bins are shown very rarely unless customers actively look for them. We see that the first five bins contain mostly high visibility brands, meaning that relevant products from low visibility brands are rarely shown to the customer unless they start to use product filters or full text search. Again, there are various reasons why a customer would prefer high visibility brands over low visibility brands, but if we wanted to (or had to because of legislation) mitigate the visibility disparities, FAIM provides a convenient approach (see Section 5.3 for a discussion on FAIM’s legal meaning for an e-commerce scenario such as ours). Figures 4c– 4f show rankings after FAIM has been applied with different values for $\theta^{A}$, $\theta^{B}$, and $\theta^{C}$. In all cases, FAIM distributes high product ranking scores more evenly across both brand visibility groups, particularly for relevant products, which is the interesting case for an e-commerce platform such as Zalando.
@@ 5.1 Recidivism Prediction
In this paper, we consider recidivism prediction for technical and argumentative reasons: the original papers by Kleinberg et al. [73] and Chouldechova [22]showing incompatibility between fairness metrics did use the COMPAS case, and we directly build on their work. Hence, we use COMPAS as an example to clarify how FAIM works. However, when considering recidivism prediction, the first questions that arise from a juridical perspective are whether one legally may, or policy-wise should, use algorithmic tools such as the COMPAS model to assess recidivism risk in Criminal Justice proceedings at all (see also 6.4). We do not wish to make any claim as to the legitimacy of the usage of these tools in this paper; there are, if anything, good reasons to be quite skeptical about it.
@@
As mentioned, COMPAS is a proprietary model used in pretrial settings in the US to determine whether potential offenders should be detained until their trial, based on a prediction of recidivism likelihood. Similar models are currently employed in Canada, the UK, and Spain [115, p. 122]. While the use of such instruments has sparked significant controversy, and critique, in the legal literature [106, 83, 101, 70, 30, 65, 97], courts such as the Wisconsin Supreme Court have condoned their use under certain safeguards,888 State v. Loomis, 881 N.W.2d 749 (Wis. 2016). and their deployment is currently on the rise [57]; [115, p. 122].
Within the scope of this paper, we cannot offer an in-depth discussion of the promises and perils of algorithmic recidivism risk assessment (see, e.g., [104, 74], and 6.4). Rather, we would like to point out that, to the extent that such tools are used at all, they must clearly fulfill minimum requirements seeking to safeguard normative and legal principles of the jurisdictions in which they are deployed. Importantly, the impossibility theorems concerning various fairness metrics mentioned above apply equally in the case of algorithmic and human risk assessments. While the influence of fairness considerations on human risk predictions, for example by a judge, remains difficult to elucidate, FAIM allows to make the relevant trade-offs transparent and to rank the involved fairness metrics in varying degrees of priority.